OpenStudy (anonymous):

What is the difference of three and seven over eight minus one and three over four in simplest form? two and one over eight two and one over two 3 2

hero (hero):

\[\left(3 + \frac{7}{8}\right) - \left(1 + \frac{3}{4}\right)\]

hero (hero):

This is what you have to begin with

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

aha

hero (hero):

Keep in mind that you can distribute the negative and remove the parentheses: \[3 + \dfrac{7}{8} - 1 -\dfrac{3}{4}\]

OpenStudy (anonymous):

yep

hero (hero):

Afterwards you can pair the whole numbers together and the fractions together

OpenStudy (anonymous):

okay

hero (hero):

\[3 - 1 + \dfrac{7}{8} - \dfrac{3}{4}\]

hero (hero):

You know what 3 - 1 is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes

hero (hero):

3 - 1 = 2 so you have \[2 + \dfrac{7}{8} - \dfrac{3}{4}\]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

:)

hero (hero):

Now try to figure out what \(\dfrac{7}{8} - \dfrac{3}{4}\) equals

OpenStudy (anonymous):

8-4=4 7-3=4 4/4 = 1 whole

hero (hero):

We don't subtract fractions like that

OpenStudy (anonymous):

okay

hero (hero):

We can only subtract fractions properly if they have the same denominator

OpenStudy (anonymous):

so i simplify

hero (hero):

You have to learn how to subtract fractions first before attempting this problem.

OpenStudy (anonymous):

hello @WHAT?!

OpenStudy (anonymous):

31/8 − 7/4

OpenStudy (anonymous):

YES THAT IS WHAT I DID

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

31/8 − 7/4 = ((31 × 4) − (7 × 8)) / (8 × 4) = (124 - 56) / 32 = 68/32 = 17/8 = 2 1/8

OpenStudy (anonymous):

okay

hero (hero):

@carolinagb02, you know how to do that but you don't know how to do \(\dfrac{7}{8} - \dfrac{3}{4}\)

OpenStudy (anonymous):

@WHAT?!

OpenStudy (anonymous):

ARE YOU THERE?

OpenStudy (anonymous):

7/8 − 3/4 = ((7 × 4) − (3 × 8)) / (8 × 4) = (28 - 24) / 32 = 4/32 = 1/8

OpenStudy (anonymous):

@Hero , please don't do that

OpenStudy (anonymous):

just \[\frac{Log4}{Log8}\]

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

@angel12310 what does that mean?

OpenStudy (anonymous):

what does what mean?

OpenStudy (anonymous):

what does the list mean?

hero (hero):

@WHAT?! please do not post your question in someone else's question

OpenStudy (anonymous):

okay guys, lets make something clear, I AM 11 YRS OLD NOT 22!

OpenStudy (anonymous):

i don't understand certain things

OpenStudy (anonymous):

plz guys don't leave!

OpenStudy (anonymous):

i'm 11 too

hero (hero):

@carolinagb02, you think you need to be 22 in order to understand how to do \[\dfrac{7}{8} - \dfrac{3}{4}\]

OpenStudy (anonymous):

7/8 − 3/4 = ((7 × 4) − (3 × 8)) / (8 × 4) = (28 - 24) / 32 = 4/32 = 1/8

OpenStudy (anonymous):

No, she thinks you have to be 22 to do\[\frac{ \log4 }{ \log8 }\]

OpenStudy (anonymous):

thats wierd

OpenStudy (anonymous):

0.66666666666

OpenStudy (anonymous):

@carolinagb02 anymore questions??

hero (hero):

@angel12310 \[\frac{7}{8} - \frac{3}{4} \\= \frac{7}{8} - \frac{3}{4} \times \dfrac{2}{2} \\=\frac{7}{8} - \frac{6}{8}\\=\frac{1}{8}\]

OpenStudy (anonymous):

I just wanted 1 answer!

OpenStudy (anonymous):

but I get a bunch of tests ;( and I came here for 1 simple answer

OpenStudy (anonymous):

Well, take Hero's word for it. I'm sure he's far more educated than all of us.

hero (hero):

Of course there's only one answer, but you have to understand how to get that answer.

hero (hero):

Subtracting the fraction is only part of the solution.

OpenStudy (anonymous):

Besides, OpenStudy isn't here to do your work for you, it's to show you how to do it so you can do it yourself later, when you can't ask other people for help. Like if you're taking a test or a quiz.

OpenStudy (anonymous):

can someone give me that one answer I HAVE BEEN HERE FOR 33 MINUTES ON THE SAME QUESTION!!!!!!! it is frustrating

hero (hero):

If we put all the steps together, we get: \[\left(3 + \frac{7}{8}\right) - \left(1 + \frac{3}{4}\right)\] \[3 - 1 + \dfrac{7}{8} - \dfrac{3}{4}\] \[2 + \frac{1}{8}\]

hero (hero):

@carolinagb02 you got stuck after I asked you to subtract the fractions.

OpenStudy (anonymous):

I know but... YOU KNOW WHAT! WHATEVER I CANT FIND ONE DUMB QUESTION I QUIT OPENSTUDY AND IT IS ALL OF UR PPL'S FAULT!

hero (hero):

Don't get frustrated. Just review how to add and subtract fractions and you'll be fine.

OpenStudy (anonymous):

I JUST WANTED TO FIGURE SOMETHING OUT AND ALL I GOT WAS A TOTAL DISSAPOINTMENT!

OpenStudy (anonymous):

I quit this job

hero (hero):

@carolinagb02, it is impossible to master adding and subtracting mixed fractions until you have mastered adding and subtracting regular fractions.

OpenStudy (anonymous):

someone PLEASE! give me and answer and stop criticizing me!

hero (hero):

You cannot do this: \[\left(3 + \frac{7}{8}\right) - \left(1 + \frac{3}{4}\right)\] Before knowing how to do this: \[\frac{7}{8} - \frac{3}{4}\]

OpenStudy (anonymous):

and if u think ur not well you are! it is mean, unkind, disrespectful and annoying, P.S it shames the person!

OpenStudy (anonymous):

you see! you just highlight peoples weaknesses!

OpenStudy (anonymous):

@Hero this is why im not ur fan!

OpenStudy (anonymous):

i gave the answer to you

OpenStudy (anonymous):

31/8 − 7/4 = ((31 × 4) − (7 × 8)) / (8 × 4) = (124 - 56) / 32 = 68/32 = 17/8 = 2 1/8

OpenStudy (anonymous):

7/8 − 3/4 = ((7 × 4) − (3 × 8)) / (8 × 4) = (28 - 24) / 32 = 4/32 = 1/8

hero (hero):

@angel12310 , 2 1/8 can easily be confused with 21/8 It is not a good notation to use. 2 + 1/8 is the correct notation

OpenStudy (anonymous):

what is a,b,c,and d

OpenStudy (anonymous):

the 1st answer, the 2nd one the 3rd etc it goes by letter!

hero (hero):

@carolinagb02, I know you are frustrated, but I cannot allow you to just simply ask for the answer. That's not how OS works

OpenStudy (anonymous):

@Hero , how are you a moderator?

hero (hero):

@carolinagb02, I understand that you are frustrated. Questioning my position as moderator really is not the issue.

OpenStudy (anonymous):

he is smarter you should ask him if it's A.,B.,C.,or D.

OpenStudy (anonymous):

I know, I just don't know if your a moderator because you have a 99 on ur status, I am not being rude, I just want to know how,

OpenStudy (anonymous):

he is smart and very helpful that's why?

OpenStudy (anonymous):

okay, cool

hero (hero):

@carolinagb02, re-post the question. I will post the complete steps one last time.

OpenStudy (anonymous):

no, I have a different one, do you know wich one it is?

OpenStudy (anonymous):

ok